OpenStudy (anonymous):

based on the graph of y=x^2-2x-3 what is the x coordinate of the positive x-intercept?

OpenStudy (anonymous):

this can be factored to y = (x-3)(x+1). x-intercept --> when y = 0. y = 0 when (x-3) = 0 or when (x+1) = 0 so x = -1 and 3. the positive one is 3.

OpenStudy (anonymous):

so its x=0, x=3? @Euler271

OpenStudy (whpalmer4):

The curve crosses the x-axis at two points, with the x-values being the solutions that make y = 0. The positive x-intercept is at x = 3. The y-value there is 0, so the fully-specified coordinate would be (3,0), but the problem only asks for the x coordinate, which is simply 3.

OpenStudy (whpalmer4):